Consider the equilibrium between COBr2, CO and Br2.

  • COBr2(g) 7e50f1f7-d29b-4777-b1ee-365a063b4be8.gifCO(g) + Br2(g) K = 0.254 at 350 K

The reaction is allowed to reach equilibrium in a 6.40-L flask. At equilibrium, [COBr2] = 0.294 M, [CO] = 0.274 M and [Br2] = 0.274 M.

(a) The equilibrium mixture is transferred to a 12.8-L flask. In which direction will the reaction proceed to reach equilibrium?


(b) Calculate the new equilibrium concentrations that result when the equilibrium mixture is transferred to a 12.8-L flask.

[COBr2] = M
[CO] = M
[Br2] = M
Answer #1

The given equilibrium reaction is COBr2(g) \rightleftharpoons CO(g) + Br2(g)

a). Initially, the reaction is allowed to reach equilibrium in a 6.40-L flask. After that, the equilibrium mixture is transferred to a 12.8-L flask i.e volume becomes doubled and when there is an increase in volume, the equilibrium will shift to the direction that produces more moles of gaseous substances. In the given equilibrium, more number of gaseous substances are present in the products side. Therefore, the equilibrium will be favored in the forward direction.

b). The new equilibrium concentrations that result when the equilibrium mixture is transferred to a 12.8-L flask.

[COBr2] = 6.4*0.294/12.8 = 0.147 M

[CO] = 6.4*0.274/12.8 = 0.137 M

[Br2] = 6.4*0.274/12.8 = 0.137 M